OpenStudy (anonymous):

evaluate (2x-y)/4 for x=6 and y=5

OpenStudy (anonymous):

plug it into the equation (2x6)-5/4 the answer is 7/4

OpenStudy (mr.math):

Replace x with 6 and y with 5, you get: \[\frac{2(6)-5}{4}=\frac{12-5}{4}=\frac{7}{4}.\]

OpenStudy (anonymous):

ok cool it is clear to me and thx to both of u... :)

OpenStudy (mr.math):

You're welcome.

OpenStudy (anonymous):

no problem =)

OpenStudy (anonymous):

good answer given... lol to both